OpenStudy (decarr432):

If f(x)=-7x+2 and g(x)=square of x+3 what is (f^o g)(-2)

OpenStudy (anonymous):

lol "square of x +3 is rather ambiguous

OpenStudy (decarr432):

square root lol

OpenStudy (anonymous):

oooh

OpenStudy (anonymous):

whatever it is, do that first

OpenStudy (anonymous):

\[g(x)=\sqrt{x+3}\] what is \(g(-2)\)?

OpenStudy (decarr432):

-7(-2)+2=16

OpenStudy (anonymous):

CAREFUL HERE

OpenStudy (anonymous):

you want \(f\circ g(-2)\) right?

OpenStudy (decarr432):

yes

OpenStudy (anonymous):

that means you have to do \(g\) FIRST then \(f\) SECOND

OpenStudy (decarr432):

Okay

OpenStudy (anonymous):

i know it is confusing because we read from left to right, so it might seem like it is backwards

OpenStudy (anonymous):

but we put the variable on the right when we write \(f(x)\) so we are always reading it backwards

OpenStudy (anonymous):

in other words \[f\circ g(-2)\] means \[f(g(-2))\] you do \(g\) first, \(f\) second

OpenStudy (decarr432):

so it would be 1 right

OpenStudy (decarr432):

x+3 would be -2+3=1 then square root?

OpenStudy (anonymous):

\[g(-2)=1\] right now do \(f(1)\)

OpenStudy (decarr432):

-7(1)+2=-5?

OpenStudy (anonymous):

bingo

OpenStudy (anonymous):

15 or 20 more and you will have mastered this for sure

OpenStudy (decarr432):

I bet lol thanks so much I'll try some on my own and if i need help I'll ask thanks again